# Generative Photomontage

### v2

@@ Appendix E Result Details
Figure 16: Additional Result: Little Red Riding Hood. (a) Vanilla ControlNet struggles to adhere to the long, complicated prompt, i.e., it assigns the wrong color to scene elements. (b) The user can instead generate a stack of images with multiple, short prompts, where each spatial region has at least one correct image within the stack. User strokes are shown on top of each image. (c) Generative Photomontage composites the user-selected regions together (bottom), where each scene element has the correct color according to the original, long prompt. Top: feature-space graph-cut result.
@@
Figure 17: Qualitative Comparison (Additional Baselines). Leftmost column: Image regions to blend (output of graph-cut optimization). Graph-cut in diffusion feature space is visualized on the left, and the image-space composite of that graph-cut is visualized on the right. CollageDiffusion [64] may struggle to preserve local appearance (b, c, d) or blend regions harmoniously (a, b, e, f). Deep Image Blending [83] may also fail to preserve local appearances (a) or show artifacts in the blended regions (b-f). Blended Latent Diffusion + MultiDiffusion [4, 8] may show artifacts at the seams, where the noise of overlapping regions are blended together (b, e), or fail to preserve local apperances (a, c, d, f).
@@
Figure 18: Qualitative Comparison (Additional Baselines). Leftmost column: Image regions to blend (output of graph-cut optimization). Graph-cut in diffusion feature space is visualized on the left, and the image-space composite of that graph-cut is visualized on the right. GP-GAN [79] does not preserve local appearances and tends to smooth out color. Cross-Domain Compositing [29] may struggle to blend regions harmoniously, particularly in non-oil painting style images (a-c, e-f). PCT-Net [28] changes the interior color of local regions but struggles to blend away the seams.
@@
Figure 22: Ablation: Injection Time Steps. Experiments of injecting $Q^{\text{comp}}$, $K^{\text{comp}}$, $V^{\text{comp}}$ after a number of time steps ($t_{\text{start}}$) during the denoising process (total: $20$ time steps). As shown, starting the injection later leads to more artifacts. Our results ($t_{\text{start}}=0$) completely removes the extra rock at the apple bite and also removes the shadow of the extra rock. However, when $t_{\text{start}}=5$, we see the shadow still remains on the base rock. This aligns with previous observations that earlier time steps form image layout and shape [86, 14], so starting the injection process later reduces the model’s ability to adapt image structure near the seams.

### v3

@@ Appendix I Limitations
Figure 16: Result: Prompt Alignment. (a) ControlNet input condition. (b) Vanilla ControlNet struggles to adhere to the long, complicated prompt. (c) With Generative Photomontage, the user can instead generate a stack of images with multiple, short prompts, where each spatial region has at least one correct image within the stack. Generative Photomontage composites the user-selected regions together, where each scene element has the correct color according to the original, long prompt.
Figure 17: Result: Prompt Alignment. Another example of using Generative Photomontage to increase prompt alignment. (a) Vanilla ControlNet struggles to adhere to the long, complicated prompt, i.e., it assigns the wrong color to scene elements. (b) The user can instead generate a stack of images with multiple, short prompts, where each spatial region has at least one correct image within the stack. User strokes are shown on top of each image. (c) Generative Photomontage composites the user-selected regions together (bottom), where each scene element has the correct color according to the original, long prompt. Top: feature-space graph-cut result.
Figure 18: Result: Prompt Alignment. (a) ControlNet input condition. (b) Vanilla ControlNet struggles to follow all aspects of the long prompt. (c) The user can instead break up the prompt into multiple, shorter prompts, and use our method to composite the outputs to create the desired image.
Figure 19: Result: Prompt Alignment. (a) ControlNet input condition. (b) Due to the unconventional shape and prompt combination (i.e., cave, eye shape, and wonderland), vanilla ControlNet struggles to adhere to all aspects of the prompt. (c) Using Generative Photomontage, users can generate multiple results with varying complexity of prompts, and composite the results to form the desired image.
Figure 20: Qualitative Comparison (Baselines). Leftmost column: Image regions to blend (output of graph-cut optimization). Graph-cut in diffusion feature space is visualized on the left, and the image-space composite of that graph-cut is visualized on the right. Interactive Digital Photomontage [1]: pixel-space graph-cut may cause seams to fall on undesired edges (see Figure 3 also), and their gradient-domain blending often fails to preserve color, e.g., the bird’s yellow beak is not preserved in (f). Blended latent diffusion [4] and MasaCtrl+ControlNet [15] may lead to color changes (c, f) and structure changes (a, b, d, e).
@@
Figure 21: Qualitative Comparison (Baselines). Leftmost column: Image regions to blend (output of graph-cut optimization). Graph-cut in diffusion feature space is visualized on the left, and the image-space composite of that graph-cut is visualized on the right. CollageDiffusion [69] may struggle to preserve local appearance (b, c, d) or blend regions harmoniously (a, b, e, f). Deep Image Blending [89] may also fail to preserve local appearances (a) or show artifacts in the blended regions (b-f). Blended Latent Diffusion + MultiDiffusion [4, 8] may show artifacts at the seams, where the noise of overlapping regions are blended together (b, e), or fail to preserve local appearances (a, c, d, f).
@@
Figure 22: Qualitative Comparison (Baselines). Leftmost column: Image regions to blend (output of graph-cut optimization). Graph-cut in diffusion feature space is visualized on the left, and the image-space composite of that graph-cut is visualized on the right. GP-GAN [85] does not preserve local appearances and tends to smooth out color. Cross-Domain Compositing [30] may struggle to blend regions harmoniously, particularly in non-oil painting style images (a-c, e-f). PCT-Net [29] changes the interior color of local regions but struggles to blend away the seams.
@@
Figure 26: Ablation: Injection Time Steps. Experiments of injecting $Q^{\text{comp}}$, $K^{\text{comp}}$, $V^{\text{comp}}$ after a number of time steps ($t_{\text{start}}$) during the denoising process (total: $20$ time steps). As shown, starting the injection later leads to more artifacts. Our results ($t_{\text{start}}=0$) completely remove the extra rock at the apple bite and also remove the shadow of the extra rock. However, when $t_{\text{start}}=5$, we see the shadow still remains on the base rock. This aligns with previous observations that earlier time steps form image layout and shape [92, 15], so starting the injection process later reduces the model’s ability to adapt image structure near the seams.
Figure 29: Generative Photomontage with SDXL. Examples of using Generative Photomontage on SDXL.
Appendix K Changelog
v1: Original draft.
v2: Fixed typos.
v3: Updated with CVPR 2025 camera ready version. Updated user survey with more responses.